### v1

# Neutrino observables in gauged $U(1)_{L_{\alpha}-L_{\beta}}$ models with two Higgs doublet and one singlet scalars

Abstract
We discuss neutrino sector in models with two Higgs doublet and one singlet scalar fields under local $U(1)_{L_{\alpha}-L_{\beta}}$ symmetry. A neutrino mass matrix is formulated for these models where the matrix is generated via type-I seesaw mechanism introducing right-handed neutrinos. The neutrino mass matrix has more degrees of freedom compared to minimal scenarios which have only one new scalar field, but its structure is still restricted by the symmetry. Then it is find that sum of neutrino mass can be lower than minimal scenarios and it is easier to satisfy observed constraints. In addition, we can fit neutrino data for $U(1)_{L_{e}-L_{\mu(\tau)}}$ cases which are disfavored in minimal models. Furthermore, some correlations among sum of neutrino mass and CP violating phases are still found although we have more free parameters.
I Introduction
The nature of neutrino such as non-zero masses and mixings is one of the biggest mystery in particle physics. We need to extend the standard model (SM) to generate neutrino masses where an attractive way is the type-I seesaw mechanism introducing right-handed neutrinos Minkowski:1977sc ; Yanagida:1979as ; Yanagida:1979gs ; Gell-Mann:1979vob ; Mohapatra:1979ia ; Schechter:1980gr as a simple possibility. New physics sector would have more rich structure containing new particle contents in addition to right-handed neutrinos and/or new (gauge) symmetries. It is thus important to explore compatibility among neutrino mass/mixing and new physics, especially a flavor dependent symmetry.
An introduction of new $U(1)$ gauge symmetry provides simple extensions of the SM. Among many possibilities local $U(1)_{L_{\alpha}-L_{\beta}}$ symmetries are interesting where they are anomaly free and the SM leptons have flavor dependent charge $L_{\alpha}-L_{\beta}(\alpha,\beta=\{e,\mu,\tau\})$. Introduction of such a symmetry restricts the structure of Yukawa interactions and Majorana mass of right-handed neutrinos that are related to neutrino mass structure. Actually compatibility among the $U(1)_{L_{\alpha}-L_{\beta}}$ symmetry and neutrino masses/mixings has been explored for type-I seesaw scenario Asai:2018ocx ; Asai:2017ryy where one scalar field, $SU(2)_{L}$ singlet or doublet with non-zero $U(1)_{L_{\alpha}-L_{\beta}}$ charge, is introduced to develop a vacuum expectation value (VEVs) contributing to neutrino mass via Yukawa interaction as minimal choices  [^1] One can find other approaches, especially $U(1)_{L_{\mu}-L_{\tau}}$ case, discussing constrained neutrino mass matrix in other seesaw models including radiative mass generation Baek:2015mna ; Lee:2017ekw ; Chen:2017gvf ; Nomura:2018vfz ; Nomura:2018cle ; Araki:2019rmw ; Asai:2019ciz ; Bauer:2020itv ; Nomura:2023vmh . . In this case we have clear relations among neutrino observables and models are restricted by the current neutrino data Esteban:2020cvm . It is then found that only $U(1)_{L_{\mu}-L_{\tau}}$ case with one singlet scalar having charge ±$\pm 1$ can accommodate neutrino data when we consider constraints on sum of neutrino mass $\sum m_{\nu}\lesssim 0.12$ eV from CMB data by Planck Planck:2018vyg under the standard $\Lambda$CDM cosmological model. Remarkably more stringent constraint is obtained as $\sum m_{\nu}<0.07$ eV (95%) if we include recent baryon acoustic oscillation (BAO) analysis by Dark Energy Spectroscopic Instrument (DESI) data DESI:2024mwx ; even stronger bounds are estimated in refs. Craig:2024tky ; Wang:2024hen . Then $U(1)_{L_{\alpha}-L_{\beta}}$ models are excluded when we only add one new scalar field and impose the constraint on the sum; it is marginal even if we consider $\sum m_{\nu}>0.059$ eV prior leading bit looser constraint $\sum m_{\nu}<0.113$. These analysis indicate we need to extend minimal $U(1)_{L_{\alpha}-L_{\beta}}$ models or modify standard $\Lambda$CDM model. Thus it is worth considering a next minimal case including both $SU(2)_{L}$ singlet and doublet scalar fields  [^2] Such a hybrid type is discussed in ref. Asai:2024pzx to relax constraints regarding $Z^{\prime}$ boson. and investigate if we can still get some predictions in neutrino sector satisfying the constraint on neutrino mass at the same time.
In this paper, we discuss models with a local $U(1)_{L_{\alpha}-L_{\beta}}$ symmetry introducing right-handed neutrinos, and $SU(2)_{L}$ doublet and singlet scalars with non-zero charges under the new gauge symmetry. The models are characterized by the gauge symmetry and possible choices of charges for doublet scalar field. We then explore neutrino masses and mixings in each model, and search for parameters that can accommodate neutrino data. For allowed parameter sets we show some predictions such as sum of neutrino mass and CP violating phases.
This paper is organized as follows. In Sec. II, we introduce models and formulate scalar sector, neutral gauge bosons, charged lepton mass and neutrino mass matrix. In Sec. III, we show numerical analysis of neutrino mass showing some predictions. Finally we conclude and discuss in Sec.IV.
II Models
 $L_{L_{i}}$  $e_{R_{i}}$  $N_{R_{i}}$  $H_{1}$  $H_{2}$  $\varphi$
$SU(2)_{L}$ $\bm{2}$ $\bm{1}$ $\bm{1}$ $\bm{2}$ $\bm{2}$ $\bm{1}$
$U(1)_{Y}$ $-\frac{1}{2}$ $-1$ 00 $1/2$ $1/2$ 00
$U(1)_{L_{\alpha}-L_{\beta}}$ $L_{\alpha}-L_{\beta}$ $L_{\alpha}-L_{\beta}$ $L_{\alpha}-L_{\beta}$ $Q_{H}$ 00 $1$
Table 1: Charge assignments of the leptons and scalar fields under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{L_{\alpha}-L_{\beta}}$ where $L_{\alpha,\beta}$ corresponds to lepton number for $\alpha(\beta)=\{e,\mu,\tau\}$ flavor ($\alpha\neq\beta$) and $Q_{H}=+1$ or $-1$.
In this section we introduce models and formulate neutrino mass matrix. Models are constructed under a framework of two Higgs doublet plus one singlet scalar under $U(1)_{L_{\alpha}-L_{\beta}}$ gauge symmetry. We also introduce three right-handed neutrinos charged under $U(1)_{L_{\alpha}-L_{\beta}}$ to realize type-I seesaw mechanism. In this scenario one Higgs doublet $H_{1}$ has charge $Q_{H}=\pm 1$ and singlet scalar has charge $1$ under $U(1)_{L_{\alpha}-L_{\beta}}$ where we summarized charge assignment in Table 1. Notice that $Q_{H}=\pm 1$ is also required to make operator $\varphi H_{1}^{\dagger}H_{2}$(or $\varphi^{*}H_{1}^{\dagger}H_{2}$) gauge invariant to avoid massless Goldstone boson from Higgs doublets. In fact these assignments of $U(1)_{L_{\alpha}-L_{\beta}}$ charge to scalar fields are only relevant ones to obtain neutrino mass matrix which can fit the neutrino data.
Relevant Lagrangian for lepton sector is written by
$L_{\ell}=$ $\ y_{\ell_{1}}\bar{L}_{L}e_{R}H_{1}+y_{\ell_{2}}\bar{L}_{L}e_{R}H_{2}+y_{\nu_{1}}\bar{L}_{L}\nu_{R}\tilde{H}_{1}+y_{\nu_{2}}\bar{L}_{L}\nu_{R}\tilde{H}_{2}$
$+\frac{1}{2}M_{0}\overline{\nu^{c}_{R}}\nu_{R}+\frac{1}{2}y_{M}\overline{\nu^{c}_{R}}\nu_{R}\varphi+\frac{1}{2}\tilde{y}_{M}\overline{\nu^{c}_{R}}\nu_{R}\varphi^{*}+h.c.\,,$ (1)
where $\tilde{H}_{i}=H_{i}^{*}i\sigma_{2}$ with $\sigma_{2}$ being the second Pauli matrix and flavor index is omitted. The structure of Yukawa coupling matrices and bare Majorana mass matrix are constrained by charge assignments under $U(1)_{L_{\alpha}-L_{\beta}}$ as we discuss below. Scalar potential is also given by
$V=$ $\ m_{1}^{2}H_{1}^{\dagger}H_{1}+m_{2}^{2}H_{2}^{\dagger}H_{2}+m_{\varphi}^{2}|\varphi|^{2}-(\mu\varphi^{(*)}H_{1}^{\dagger}H_{2}+h.c.)+\lambda_{1}(H_{1}^{\dagger}H_{1})^{2}+\lambda_{2}(H_{2}^{\dagger}H_{2})^{2}+\lambda_{\varphi}|\varphi|^{4}$
$+\lambda_{3}(H_{1}^{\dagger}H_{1})(H_{2}^{\dagger}H_{2})+\lambda_{4}(H_{1}^{\dagger}H_{2})(H_{2}^{\dagger}H_{1})+\lambda_{H_{1}\varphi}(H_{1}^{\dagger}H_{1})|\varphi|^{2}+\lambda_{H_{2}\varphi}(H_{2}^{\dagger}H_{2})|\varphi|^{2},$ (2)
where $\varphi$ or $\varphi^{*}$ is determined by the $U(1)_{L_{\alpha}-L_{\beta}}$ charge of $H_{1}$ for the fourth term. In this work we consider six models given in Table 2 that are distinguished by $U(1)$ symmetry and charge $Q_{H}$ providing us different structure of neutrino mass matrix.
 model (1)  model (2)  model (3)  model (4)  model (5)  model (6)
$U(1)_{L_{i}-L_{j}}$ $L_{e}-L_{\mu}$ $L_{e}-L_{\mu}$ $L_{e}-L_{\tau}$ $L_{e}-L_{\tau}$ $L_{\mu}-L_{\tau}$ $L_{\mu}-L_{\tau}$
$Q_{H}$ $1$ $-1$ $1$ $-1$ $1$ $-1$
Table 2: Models distinguished by extra $U(1)$ symmetry and $H_{1}$ charge $Q_{H}$.
II.1 Scalar sector
Here we review scalar sector containing two Higgs doublet and one singlet scalar under extra $U(1)_{L_{\alpha}-L_{\beta}}$ gauge symmetry. The two Higgs doublets and singlet scalar are represented as
$H_{i}=\begin{pmatrix}\phi_{a}^{+}\\ \frac{1}{\sqrt{2}}(v_{a}+h_{a}+i\eta_{a})\end{pmatrix},\quad\varphi=\frac{1}{\sqrt{2}}(\phi_{R}+v_{\varphi}+i\phi_{I}),$ (3)
where $a=1,2$, and $v_{a}$ and $v_{\varphi}$ are the VEVs of the corresponding fields. The VEVs can be obtained from the stationary conditions $(\partial V/\partial h_{i})_{0}=(\partial V/\partial\phi_{R})_{0}=0$ where subscript "$0"$ indicates all the component fields are taken to be zero. These conditions provide
$m_{H_{1}}^{2}v_{1}-m_{12}^{2}v_{2}+\frac{v_{1}}{2}(2v_{1}^{2}\lambda_{1}+v_{2}^{2}\bar{\lambda}+v_{\varphi}^{2}\lambda_{H_{1}\varphi})=0$
$m_{H_{2}}^{2}v_{2}-m_{12}^{2}v_{1}+\frac{v_{2}}{2}(2v_{2}^{2}\lambda_{2}+v_{1}^{2}\bar{\lambda}+v_{\varphi}^{2}\lambda_{H_{2}\varphi})=0$
$m_{\varphi}^{2}v_{\varphi}-\frac{1}{\sqrt{2}}\mu v_{1}v_{2}+\lambda_{\varphi}v_{\varphi}^{3}+\frac{\lambda_{H_{1}\varphi}}{2}v_{1}^{2}v_{\varphi}+\frac{\lambda_{H_{2}\varphi}}{2}v_{2}^{2}v_{\varphi}=0,$ (4)
where $m_{12}^{2}\equiv\mu v_{\varphi}/\sqrt{2}$ and $\bar{\lambda}=\lambda_{3}+\lambda_{4}$.
The mass eigenstates for charged scalar components are obtained as in a two Higgs doublet model (THDM),
$\begin{pmatrix}G^{\pm}\\ H^{\pm}\end{pmatrix}=\begin{pmatrix}\cos\beta&-\sin\beta\\ \sin\beta&\cos\beta\end{pmatrix}\begin{pmatrix}\phi_{1}^{\pm}\\ \phi_{2}^{\pm}\end{pmatrix},$ (5)
where $\tan\beta=v_{2}/v_{1}$, $G^{\pm}$ corresponds to Nambu-Goldstone (NG) boson absorbed by $W^{\pm}$ and $H^{\pm}$ is physical charged Higgs boson. The mass of the charged Higgs boson can be written by
$m^{2}_{H^{\pm}}=\frac{m^{2}_{12}}{\sin\beta\cos\beta}-\frac{v^{2}}{2}\lambda_{4},$ (6)
where $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\simeq 246$ GeV.
Applying conditions in Eq. (4), the mass matrix for CP-odd scalar bosons is obtained as
$\mathcal{L}\in\frac{1}{2}\begin{pmatrix}\eta_{1}\\ \eta_{2}\\ \phi_{I}\end{pmatrix}^{T}\begin{pmatrix}\frac{\mu v_{\varphi}v_{2}}{\sqrt{2}v_{1}}&-\frac{\mu v_{\varphi}}{\sqrt{2}}&-\frac{\mu v_{2}}{\sqrt{2}}\\ -\frac{\mu v_{\varphi}}{\sqrt{2}}&\frac{\mu v_{\varphi}v_{1}}{\sqrt{2}v_{2}}&\frac{\mu v_{1}v_{2}}{\sqrt{2}v_{\varphi}}\\ -\frac{\mu v_{2}}{\sqrt{2}}&\frac{\mu v_{1}}{\sqrt{2}}&\frac{\mu v_{2}}{\sqrt{2}}\end{pmatrix}\begin{pmatrix}\eta_{1}\\ \eta_{2}\\ \phi_{I}\end{pmatrix}.$ (7)
This mass matrix can be diagonalized by rotating the basis as follows Nomura:2019wlo ;
$\begin{pmatrix}\eta_{1}\\ \eta_{2}\\ \phi_{I}\end{pmatrix}=\begin{pmatrix}\frac{v_{1}}{v}&-\frac{v_{\varphi}v_{2}}{\sqrt{v_{1}^{2}v_{2}^{2}+v_{\varphi}^{2}v^{2}}}&\frac{v_{1}}{\sqrt{v_{\varphi}^{2}+v_{1}^{2}}}\\ \frac{v_{2}}{v}&-\frac{v_{\varphi}v_{1}}{\sqrt{v_{1}^{2}v_{2}^{2}+v_{\varphi}^{2}v^{2}}}&0\\ 0&-\frac{v_{1}v_{2}}{\sqrt{v_{1}^{2}v_{2}^{2}+v_{\varphi}^{2}v^{2}}}&\frac{v_{\varphi}}{\sqrt{v_{\varphi}^{2}+v_{1}^{2}}}\end{pmatrix}\begin{pmatrix}G^{0}_{1}\\ A^{0}\\ G^{0}_{2}\end{pmatrix},$ (8)
where $G_{1,2}^{0}$ are massless NG boson whose degrees of freedom are absorbed by $Z$ and $Z^{\prime}$ bosons. Here we note that $\{\eta_{1},\eta_{2}\}$ sector becomes THDM like in the limit of $v_{\varphi}\gg v$. The mass eigenvalue of physical CP-odd scalar boson $A^{0}$ is given by
$m^{2}_{A^{0}}=\frac{m_{12}^{2}}{\sin\beta\cos\beta}+\frac{1}{\sqrt{2}}\frac{\mu v^{2}}{v_{\varphi}}.$ (9)
Note that the mass of $A^{0}$ becomes zero when we take $\mu=0$ since there is spontaneously broken global $U(1)$ symmetry in the potential for the limit.
The CP-even scalar sector has three physical degrees of freedom $\{h_{1},h_{2},\phi_{R}\}$ and the mass matrix is written by
$\mathcal{L}\supset\frac{1}{2}\left(\begin{array}[]{c}h_{1}\\ h_{2}\\ \phi_{R}\end{array}\right)^{T}\left(\begin{array}[]{ccc}2\lambda_{1}v_{1}^{2}+\frac{\mu v_{\varphi}v_{2}}{\sqrt{2}v_{1}}&\lambda_{3}v_{1}v_{2}+\lambda_{4}v_{1}v_{2}-\frac{\mu v_{\varphi}}{\sqrt{2}}&\eta\lambda_{H_{1}\varphi}v_{1}-\frac{\mu v_{2}}{\sqrt{2}}\\ \lambda_{3}v_{1}v_{2}+\lambda_{4}v_{1}v_{2}-\frac{\mu v_{\varphi}}{\sqrt{2}}&2\lambda_{2}v_{2}^{2}+\frac{\mu v_{1}v_{\varphi}}{\sqrt{2}v_{2}}&v_{\varphi}\lambda_{H_{2}\varphi}v_{2}-\frac{\mu v_{1}}{\sqrt{2}}\\ v_{\varphi}\lambda_{H_{1}\varphi}v_{1}-\frac{\mu v_{2}}{\sqrt{2}}&v_{\varphi}\lambda_{H_{2}\varphi}v_{2}-\frac{\mu v_{1}}{\sqrt{2}}&2v_{\varphi}^{2}\lambda_{\varphi}+\frac{\mu v_{1}v_{2}}{\sqrt{2}\varphi}\\ \end{array}\right)\left(\begin{array}[]{c}h_{1}\\ h_{2}\\ \phi_{R}\end{array}\right),$ (10)
where we imposed conditions in Eq. (4). This mass matrix can be diagonalized by $3\times 3$ orthogonal matrix providing three physical mass eigenvalues. Such an orthogonal matrix $R$ with three Euler parameters $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ is written by
$\ R(\alpha_{1},\alpha_{2},\alpha_{3})=\left(\begin{array}[]{ccc}c_{\alpha_{1}}c_{\alpha_{2}}&-s_{\alpha_{1}}c_{\alpha_{2}}&s_{\alpha_{2}}\\ -c_{\alpha_{1}}s_{\alpha_{2}}s_{\alpha_{3}}+s_{\alpha_{1}}c_{\alpha_{3}}&c_{\alpha_{1}}c_{\alpha_{3}}+s_{\alpha_{1}}s_{\alpha_{2}}s_{\alpha_{3}}&c_{\alpha_{2}}s_{\alpha_{3}}\\ -c_{\alpha_{1}}s_{\alpha_{2}}c_{\alpha_{3}}-s_{\alpha_{1}}s_{\alpha_{3}}&-c_{\alpha_{1}}s_{\alpha_{3}}+s_{\alpha_{1}}s_{\alpha_{2}}c_{\alpha_{3}}&c_{\alpha_{2}}c_{\alpha_{3}}\\ \end{array}\right),$ (11)
where $c_{\alpha_{i}}=\cos\alpha_{i}(s_{\alpha_{i}}=\sin\alpha_{i})$. Then mass eigenstates are obtained such that
$\left(\begin{array}[]{c}h_{1}\\ h_{2}\\ \phi_{R}\end{array}\right)=R_{ij}\left(\begin{array}[]{c}H^{0}\\ h^{0}\\ \xi^{0}\end{array}\right)_{j}.$ (12)
In this work, we do not discuss more details of the scalar sector since our focus is neutrino sector. Thus we simply assume parameters in the scalar potential are chosen to satisfy phenomenological constraints regarding Higgs boson physics.
II.2 Neutral gauge bosons
Here we focus on neutral gauge boson sector in the model since the charged gauge boson $W^{\pm}$ is the same as the SM one. After spontaneous symmetry breaking we obtain mass terms
$\mathcal{L}_{M}=\frac{1}{2}m^{2}_{Z_{\rm SM}}\tilde{Z}_{\mu}\tilde{Z}^{\mu}+\frac{1}{2}\tilde{m}_{Z^{\prime}}^{2}\tilde{Z}^{\prime}_{\mu}\tilde{Z}^{\prime\mu}+\Delta M^{2}\tilde{Z}_{\mu}\tilde{Z}^{\prime\mu},$ (13)
where $m^{2}_{Z_{\rm SM}}=v^{2}(g_{1}^{2}+g_{2}^{2})/4$ with $g_{1(2)}$ being $U(1)_{Y}(SU(2)_{L})$ gauge coupling, $\tilde{m}_{Z^{\prime}}^{2}=g_{X}^{2}(v_{\varphi}^{2}+v_{1}^{2})$ with new gauge coupling $g_{X}$, and $\Delta M^{2}=g_{X}\sqrt{g_{1}^{2}+g_{2}^{2}}v_{1}^{2}/2$. Here $\tilde{Z}_{\mu}=\cos\theta_{W}W^{3}_{\mu}-\sin\theta_{W}B_{\mu}$ where $W^{3}_{\mu}$ and $B_{\mu}$ come from $SU(2)_{L}$ and $U(1)_{Y}$ gauge fields as in the SM while $\tilde{Z}^{\prime}_{\mu}$ is $U(1)_{L_{\alpha}-L_{\beta}}$ gauge field. Notice that we do not consider kinetic mixing among $U(1)$ gauge fields assuming it is negligibly small. Then we can diagonalize the mass terms by rotating the fields and get mass eigenstates
$\begin{pmatrix}Z\\ Z^{\prime}\end{pmatrix}=\begin{pmatrix}\cos\chi&\sin\chi\\ -\sin\chi&\cos\chi\end{pmatrix}\begin{pmatrix}\tilde{Z}\\ \tilde{Z}^{\prime}\end{pmatrix},$ (14)
$\tan 2\chi=\frac{2\Delta M^{2}}{m^{2}_{Z_{\rm SM}}-\tilde{m}^{2}_{Z^{\prime}}}.$ (15)
Mass eigenvalues are also obtained as
$m^{2}_{Z,Z^{\prime}}=\frac{1}{2}(m^{2}_{Z_{\rm SM}}+\tilde{m}_{Z^{\prime}}^{2})\pm\frac{1}{2}\sqrt{(m^{2}_{Z_{\rm SM}}-\tilde{m}_{Z^{\prime}}^{2})^{2}+4\Delta M^{4}}.$ (16)
In this work we are not going to discuss $Z^{\prime}$ physics assuming $Z$-$Z^{\prime}$ mixing is negligibly small to avoid electroweak precision tests, and require $Z^{\prime}$ mass and gauge coupling $g_{X}$ satisfying current constraints from $Z^{\prime}$ searches.
II.3 Charged lepton mass
Mass matrix of charged lepton is generated via Yukawa interactions in Eq. (1) after electroweak symmetry breaking. Thus the mass term is given by
$\overline{\ell_{L}}(M_{e})_{\ell\ell^{\prime}}\ell^{\prime}_{R}+h.c.=\overline{\ell_{L}}\left(\frac{v_{1}}{\sqrt{2}}y_{\ell_{1}}+\frac{v_{2}}{\sqrt{2}}y_{\ell_{2}}\right)_{\ell\ell^{\prime}}\ell^{\prime}_{R}+h.c.\,,$ (17)
where ()=$\ell(\ell^{\prime})=\{e,\mu,\tau\}$. In general the mass matrix is not diagonal and its structure depends on the model. The possible structures of the matrix $M_{e}$ are as follows
$M^{\rm Model\,(1)}_{e}:\left(\begin{array}[]{ccc}\times&0&\times\\ 0&\times&0\\ 0&\times&\times\end{array}\right),\quad M^{\rm Model\,(2)}_{e}:\left(\begin{array}[]{ccc}\times&0&0\\ 0&\times&\times\\ \times&0&\times\end{array}\right),\quad M^{\rm Model\,(3)}_{e}:\left(\begin{array}[]{ccc}\times&\times&0\\ 0&\times&\times\\ 0&0&\times\end{array}\right),$ (27)
$M^{\rm Model\,(4)}_{e}:\left(\begin{array}[]{ccc}\times&0&0\\ \times&\times&0\\ 0&\times&\times\end{array}\right),\quad M^{\rm Model\,(5)}_{e}:\left(\begin{array}[]{ccc}\times&0&\times\\ \times&\times&0\\ 0&0&\times\end{array}\right),\quad M^{\rm Model\,(6)}_{e}:\left(\begin{array}[]{ccc}\times&\times&0\\ 0&\times&0\\ \times&0&\times\end{array}\right),$ (37)
where $\times$ indicates a non-zero element. The mass matrix can be diagonalized by bi-unitary matrices $V_{eL}$ and $V_{eR}$ transforming $\ell_{L}$ and $\ell_{R}$; $M_{d}^{\rm diagonal}=V_{eL}M_{e}V_{eR}^{\dagger}$. Note that such transformation induces FCNC for interaction among $Z^{\prime}$ and leptons due to lepton flavor dependent charges and it is strongly constrained by LFV search experiments Asai:2018ocx . We then assume off-diagonal elements of the matrix is negligibly small and $V_{eL(eR)}\simeq 1$ for simplicity.
II.4 Neutrino mass
After spontaneous symmetry breaking we obtain Dirac mass $M_{D}$ between $\nu_{L}$ and $\nu_{R}$ as well as Majorana mass $M_{R}$ of $\nu_{R}$. These mass matrices are written by
$M_{D}$ $=\frac{y_{\nu_{1}}v_{1}}{\sqrt{2}}+\frac{y_{\nu_{2}}v_{2}}{\sqrt{2}},$ (38)
$M_{R}$ $=M_{0}+\frac{y_{M}v_{\varphi}}{\sqrt{2}}+\frac{\tilde{y}_{M}v_{\varphi}}{\sqrt{2}}.$ (39)
Structure of these matrices is different for each model. For $M_{D}$ we find structure as follows:
$M^{\rm Model\,(1)}_{D}:\left(\begin{array}[]{ccc}\times&0&0\\ 0&\times&\times\\ \times&0&\times\end{array}\right),\quad M^{\rm Model\,(2)}_{D}:\left(\begin{array}[]{ccc}\times&0&\times\\ 0&\times&0\\ 0&\times&\times\end{array}\right),\quad M^{\rm Model\,(3)}_{D}:\left(\begin{array}[]{ccc}\times&0&0\\ \times&\times&0\\ 0&\times&\times\end{array}\right),$ (49)
$M^{\rm Model\,(4)}_{D}:\left(\begin{array}[]{ccc}\times&\times&0\\ 0&\times&\times\\ 0&0&\times\end{array}\right),\quad M^{\rm Model\,(5)}_{D}:\left(\begin{array}[]{ccc}\times&\times&0\\ 0&\times&0\\ \times&0&\times\end{array}\right),\quad M^{\rm Model\,(6)}_{D}:\left(\begin{array}[]{ccc}\times&0&\times\\ \times&\times&0\\ 0&0&\times\end{array}\right).$ (59)
We also find structure of $M_{R}$ for each model such that
$M^{\rm Model\,(1,2)}_{R}:\left(\begin{array}[]{ccc}0&\times&\times\\ \times&0&\times\\ \times&\times&\times\end{array}\right),\quad M^{\rm Model\,(3,4)}_{R}:\left(\begin{array}[]{ccc}0&\times&\times\\ \times&\times&\times\\ \times&\times&0\end{array}\right),\quad M^{\rm Model\,(5,6)}_{R}:\left(\begin{array}[]{ccc}\times&\times&\times\\ \times&0&\times\\ \times&\times&0\end{array}\right),$ (69)
where the structure is the same when a gauge symmetry of models is identical.
The neutrino mass in type-I seesaw model is given by
$m_{\nu}=-M_{D}M_{R}^{-1}M_{D}^{T}.$ (70)
We then write active neutrino mass matrix by
$m_{\nu}=\kappa\tilde{m}_{\nu},$ (71)
where $\kappa$ has mass dimension and $\tilde{m}_{\nu}$ is dimensionless. The neutrino mass matrix $m_{\nu}$ is diagonalized by a unitary matrix $V_{\nu}$ by $D_{\nu}=|\kappa|\tilde{D}_{\nu}=V_{\nu}^{T}m_{\nu}V_{\nu}=|\kappa|V_{\nu}^{T}\tilde{m}_{\nu}V_{\nu}$. Then $|\kappa|$ is determined by
$(\mathrm{NO}):\ |\kappa|^{2}=\frac{|\Delta m_{\rm atm}^{2}|}{\tilde{D}_{\nu_{3}}^{2}-\tilde{D}_{\nu_{1}}^{2}},\quad(\mathrm{IO}):\ |\kappa|^{2}=\frac{|\Delta m_{\rm atm}^{2}|}{\tilde{D}_{\nu_{2}}^{2}-\tilde{D}_{\nu_{3}}^{2}},$ (72)
where $\Delta m_{\rm atm}^{2}$ is atmospheric neutrino mass-squared splitting, and NO and IO respectively represent the normal ordering and the inverted ordering of neutrino mass eigenvalues. Subsequently, the solar mass squared splitting can be written in terms of $|\kappa|$ as follows:
$\Delta m_{\rm sol}^{2}=|\kappa|^{2}({\tilde{D}_{\nu_{2}}^{2}-\tilde{D}_{\nu_{1}}^{2}}),$ (73)
which can be compared to the observed value. The observed mixing matrix is defined by $U=V^{\dagger}_{eL}V_{\nu}$  [^3] We can approximate $V_{eL}\simeq 1$ to suppress LFV as discussed in previous subsection. , where it is parametrized by three mixing angles $\theta_{ij}(i,j=1,2,3;i<j)$, one CP violating Dirac phase $\delta_{CP}$, and two Majorana phases $\alpha_{21},\alpha_{31}$ as follows:
$U=\begin{pmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}e^{-i\delta_{CP}}\\ -s_{12}c_{23}-c_{12}s_{23}s_{13}e^{i\delta_{CP}}&c_{12}c_{23}-s_{12}s_{23}s_{13}e^{i\delta_{CP}}&s_{23}c_{13}\\ s_{12}s_{23}-c_{12}c_{23}s_{13}e^{i\delta_{CP}}&-c_{12}s_{23}-s_{12}c_{23}s_{13}e^{i\delta_{CP}}&c_{23}c_{13}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&e^{i\frac{\alpha_{21}}{2}}&0\\ 0&0&e^{i\frac{\alpha_{31}}{2}}\end{pmatrix}.$ (74)
Here, $c_{ij}$ and $s_{ij}$ stand for $\cos\theta_{ij}$ and $\sin\theta_{ij}$ ($i,j=1-3$), respectively. The matrix $U$ is identified as the PMNS matrix Pontecorvo:1957qd ; Maki:1962mu . Then, each of the mixings is given in terms of the component of $U$ as follows:
$\sin^{2}\theta_{13}=|U_{e3}|^{2},\quad\sin^{2}\theta_{23}=\frac{|U_{\mu 3}|^{2}}{1-|U_{e3}|^{2}},\quad\sin^{2}\theta_{12}=\frac{|U_{e2}|^{2}}{1-|U_{e3}|^{2}}.$ (75)
The Dirac phase $\delta_{CP}$ is given by computing the Jarlskog invariant as follows:
$\sin\delta_{CP}$ $=\frac{\text{Im}[U_{e1}U_{\mu 2}U_{e2}^{*}U_{\mu 1}^{*}]}{s_{23}c_{23}s_{12}c_{12}s_{13}c^{2}_{13}},\quad\cos\delta_{CP}=-\frac{|U_{\tau 1}|^{2}-s^{2}_{12}s^{2}_{23}-c^{2}_{12}c^{2}_{23}s^{2}_{13}}{2c_{12}s_{12}c_{23}s_{23}s_{13}},$ (76)
where $\delta_{CP}$ be subtracted from $\pi$ if $\cos\delta_{CP}$ is negative. Majorana phase $\alpha_{21},\ \alpha_{31}$ are found as
$\sin\left(\frac{\alpha_{21}}{2}\right)=\frac{\text{Im}[U^{*}_{e1}U_{e2}]}{c_{12}s_{12}c_{13}^{2}},\quad\cos\left(\frac{\alpha_{21}}{2}\right)=\frac{\text{Re}[U^{*}_{e1}U_{e2}]}{c_{12}s_{12}c_{13}^{2}},\ ,$ (77)
$\sin\left(\frac{\alpha_{31}}{2}-\delta_{CP}\right)=\frac{\text{Im}[U^{*}_{e1}U_{e3}]}{c_{12}s_{13}c_{13}},\quad\cos\left(\frac{\alpha_{31}}{2}-\delta_{CP}\right)=\frac{\text{Re}[U^{*}_{e1}U_{e3}]}{c_{12}s_{13}c_{13}},$ (78)
where $\alpha_{21}/2,\ \alpha_{31}/2-\delta_{CP}$ are subtracted from $\pi$, when $\cos\left(\frac{\alpha_{21}}{2}\right),\ \cos\left(\frac{\alpha_{31}}{2}-\delta_{CP}\right)$ are negative. In addition, the effective mass for the neutrinoless double beta decay is given by
$\langle m_{ee}\rangle=|\kappa||\tilde{D}_{\nu_{1}}\cos^{2}\theta_{12}\cos^{2}\theta_{13}+\tilde{D}_{\nu_{2}}\sin^{2}\theta_{12}\cos^{2}\theta_{13}e^{i\alpha_{21}}+\tilde{D}_{\nu_{3}}\sin^{2}\theta_{13}e^{-2i\delta_{CP}}|,$ (79)
where its observed value could be tested in experiments such as KamLAND-Zen KamLAND-Zen:2022tow , LEGEND LEGEND:2017cdu and nEXO nEXO:2017nam .
III Numerical analysis
Figure 1: The predicted values on $\{m_{1},\langle m_{ee}\rangle\}$ plane for allowed parameter points in each model.
Figure 2: The predicted values on $\{\sum m_{\nu},\delta_{\rm CP}\}$ plane for allowed parameter points in each model. The light-blue(cyan) region indicates 1(3)$\sigma$ range of $\delta_{\rm CP}$ value.
Figure 3: The predicted values on $\{\delta_{\rm CP},\alpha_{21}\}$ plane for allowed parameter points in each model. The light-blue(cyan) region also indicates 1(3)$\sigma$ range of $\delta_{\rm CP}$ value.
In this section we discuss neutrino observables such as masses, mixings and CP-phases in our models. These values are numerically estimated since structures of the neutrino mass matrix are not very simple to get analytic solutions for neutrino observables. In numerical calculation we modify neutrino mass matrix as
$m_{\nu}=-\frac{(M_{D})_{11}^{2}}{(M_{R})_{kk}}\tilde{M}_{D}\tilde{M}_{R}^{-1}\tilde{M}_{D}^{T},$ (80)
where $\tilde{M}_{D}\equiv M_{D}/(M_{D})_{11}$ and $\tilde{M}_{R}=M_{R}/(M_{R})_{kk}$ with $(M_{R})_{kk}$ being a non-zero diagonal element of $M_{R}$ in Eq. (69). The scaling factor $\frac{(M_{D})_{11}^{2}}{(M_{R})_{kk}}$ is identified as $\kappa$ in Eq. (71) that will be evaluated by Eq. (72). Then we scan the element of $\tilde{M}_{D}$ and $\tilde{M}_{R}$ in the range of
$\left|(\tilde{M}_{D,R})_{ij}\right|\in[0.1,10],\quad\left((\tilde{M}_{D})_{11}=1,\quad(\tilde{M}_{R})_{kk}=1,\right)$ (81)
[…]
$0.270<\sin^{2}\theta_{12}<0.341,\quad 0.406<\sin^{2}\theta_{23}<0.620,\quad 0.02029<\sin^{2}\theta_{13}<0.02391,$
$2.428\times 10^{-3}\ {\rm eV}^{2}<|\Delta m_{\rm sol}^{2}|<2.597\times 10^{-3}\ {\rm eV}^{2},$
$6.82\times 10^{-5}\ {\rm eV}^{2}<\Delta m^{2}_{\rm sol}<8.03\times 10^{-5}\ {\rm eV}^{2}.$ (82)
[…]
In Fig. 1, we show the $\langle m_{ee}\rangle$ and $m_{1}$ (the lightest neutrino mass) for allowed parameter points in each model. The horizontal lines shows current constraint from KamLand-Zen KamLAND-Zen:2022tow and future prospect in nEXO nEXO:2017nam with energy-density functional (EDF) theory for nuclear matrix element. The distributions are almost similar in these models and some points can be tested in future neutrinoless double beta decay experiments.
Figure 4: The predicted values on $\{\delta_{\rm CP},\alpha_{31}\}$ plane for allowed parameter points in each model. The light-blue(cyan) region is the same as Fig. 3.
In fig. 2, we show $\sum m_{\nu}$ and $\delta_{\rm CP}$ for allowed parameter points in each model. The vertical line indicates the upper limit by CMB+DESI data. We find some correlation between $\sum m_{\nu}$ and $\delta_{\rm CP}$ where models (1) and (6) [(3) and (5)] provide similar behavior. As we see, they show ditinguishable behavior where one tends to provide small $\sum m_{\nu}$ around $\delta_{\rm CP}\sim 200^{\circ}$ and the other one tends to give small $\sum m_{\nu}$ around $\delta_{\rm CP}\sim 0^{\circ}(360^{\circ})$. We also indicate 1(3)$\sigma$ range of $\delta_{\rm CP}$ value by the light-blue(cyan) region. These two types of distributions could be distinguished if we have more precision for $\sum m_{\nu}$ and $\delta_{\rm CP}$ in future.
Fig. 3 and 4 show $\delta_{\rm CP}$-$\alpha_{21}$ and $\delta_{\rm CP}$-$\alpha_{31}$ values for allowed parameter points in each model. We find some correlations among phases where correlations of $\delta_{\rm CP}$-$\alpha_{21}$ in models (1) and (6) [(3) and (5)] provide similar behavior. In each case the value of $\alpha_{21}$ is concentrated around 100-250 degrees. On the other hand correlations of $\delta_{\rm CP}$-$\alpha_{31}$ are different in each model. In these plots we show 1(3)$\sigma$ range of $\delta_{\rm CP}$ value by the light-blue(cyan) region as in the Fig. 2. In addition, we also show correlation between $\alpha_{21}$ and $\alpha_{31}$ in Fig. 5 where models (1) and (6) [(3) and (5)] provide similar behavior.
Figure 5: The predicted values on $\{\alpha_{21},\alpha_{31}\}$ plane for allowed parameter points in each model.
IV Summary and discussion
We have discussed neutrino observables in $U(1)_{L_{\alpha}-L_{\beta}}$ models where we introduced second Higgs doublet and one singlet scalar fields with non-zero charges under new $U(1)$, as extensions of minimal models that have only one new scalar field. The six models are characterized by the choice of $U(1)_{L_{\alpha}-L_{\beta}}$ gauge symmetry and its charge assignment to the second Higgs doublet. We have then formulated scalar sector, neutral gauge bosons, charged lepton mass matrix and neutrino mass matrix in the models. The structure of neutrino mass matrix is still restricted by the gauge symmetry but number of free parameters is increased compared to minimal cases.
[…]
Acknowledgments
The work was also supported by the Fundamental Research Funds for the Central Universities (T. N.).

### v2

@@
 Asai:2024pzx to relax constraints regarding $Z^{\prime}$ boson.
+Also similar hybrid models are discussed in refs.
+Heeck:2011wj ; Heeck:2014qea ; Crivellin:2015lwa ; Ardu:2022zom for lepton flavor violation physics.
 and investigate if we can still get some predictions in neutrino sector satisfying the constraint on neutrino mass at the same time.
@@
 II Models
- $L_{L_{i}}$  $e_{R_{i}}$  $N_{R_{i}}$  $H_{1}$  $H_{2}$  $\varphi$
+ $L_{L_{i}}$  $e_{R_{i}}$  $\nu_{R_{i}}$  $H_{1}$  $H_{2}$  $\varphi$
 $SU(2)_{L}$ $\bm{2}$ $\bm{1}$ $\bm{1}$ $\bm{2}$ $\bm{2}$ $\bm{1}$
@@
 $m^{2}_{Z,Z^{\prime}}=\frac{1}{2}(m^{2}_{Z_{\rm SM}}+\tilde{m}_{Z^{\prime}}^{2})\pm\frac{1}{2}\sqrt{(m^{2}_{Z_{\rm SM}}-\tilde{m}_{Z^{\prime}}^{2})^{2}+4\Delta M^{4}}.$ (16)
-In this work we are not going to discuss $Z^{\prime}$ physics assuming $Z$-$Z^{\prime}$ mixing is negligibly small to avoid electroweak precision tests, and require $Z^{\prime}$ mass and gauge coupling $g_{X}$ satisfying current constraints from $Z^{\prime}$ searches.
+Here we briefly discuss constraints on $Z$-$Z^{\prime}$ mixing from electroweak precision test (EWPT).
+The mixing is typically constrained as $|\sin\chi|\lesssim 10^{-3}-10^{-4}$ when $Z^{\prime}$ boson comes from local $U(1)$ symmetry as subgroup of GUT symmetry like $SO(10)$ and $E_{6}$ Erler:2009jh .
+EWPT constraint is discussed for $U(1)_{L_{\mu}-L_{\tau}}$ case in ref.
+Heeck:2011wj where we find $-0.0008\lesssim g_{X}\sin\chi\lesssim 0.0003$.
+We apply the constraint as a reference under $m_{Z}^{2}\ll m_{Z^{\prime}}^{2}$ mass hierarchy for neutral gauge bosons.
+In our case the constraint approximately becomes
+$\frac{g_{X}^{2}m_{Z}v\cos\beta}{m_{Z^{\prime}}^{2}}\lesssim 8\times 10^{-4}.$ (17)
+In Fig.
+1, we show the region excluded by EWPT for $U(1)_{L_{\mu}-L_{\tau}}$ with $\tan\beta=5$.
+It is shown that $\mathcal{O}(1)$ gauge coupling is allowed if $Z^{\prime}$ mass satisfies 1700 GeV $\lesssim m_{Z^{\prime}}$ where the allowed region is also safe from other experimental constraints.
+We expect similar $Z$-$Z^{\prime}$ mixing constraint in the other cases, $U(1)_{L_{e}-L_{\mu}}$ and $U(1)_{L_{e}-L_{\tau}}$.
+This constraint restricts $v_{\varphi}\simeq m_{Z^{\prime}}/g_{X}$ and can affect Majorana mass of $\nu_{R}$.
+However it does not provide much impact in fitting neutrino mass since we also have Yukawa couplings that are free parameters.
+In our analysis below we consider heavy $Z^{\prime}$ and assume $Z$-$Z^{\prime}$ mixing is negligibly small to avoid EWPT.
+Figure 1: The gray region is excluded by EWPT for $U(1)_{L_{\mu}-L_{\tau}}$ case with $\tan\beta=5$.
 II.3 Charged lepton mass
@@
 where $\times$ indicates a non-zero element.
-The mass matrix can be diagonalized by bi-unitary matrices $V_{eL}$ and $V_{eR}$ transforming $\ell_{L}$ and $\ell_{R}$; $M_{d}^{\rm diagonal}=V_{eL}M_{e}V_{eR}^{\dagger}$.
+The mass matrix can be diagonalized by bi-unitary matrices $V_{eL}$ and $V_{eR}$ transforming $\ell_{L}$ and $\ell_{R}$; $M_{d}^{\rm diagonal}=V_{eL}^{\dagger}M_{e}V_{eR}$.
 Note that such transformation induces FCNC for interaction among $Z^{\prime}$ and leptons due to lepton flavor dependent charges and it is strongly constrained by LFV search experiments Asai:2018ocx .
-We then assume off-diagonal elements of the matrix is negligibly small and $V_{eL(eR)}\simeq 1$ for simplicity.
+Here we illustrate a LFV constraint in a simple case.
+Firstly $Z^{\prime}$ interaction for charged leptons is written by
+$\mathcal{L}_{\rm Z^{\prime}\ell\bar{\ell}}=g_{X}Z^{\prime}_{\mu}\left[\overline{\ell_{L}}V^{\dagger}_{eL}Q_{L_{\alpha}-L_{\beta}}V_{eL}\gamma^{\mu}\ell_{L}+\overline{\ell_{R}}V^{\dagger}_{eR}Q_{L_{\alpha}-L_{\beta}}V_{eR}\gamma^{\mu}\ell_{R}\right],$ (39)
+where $Q_{L_{\alpha-\beta}}$ indicates diagonal matrix whose elements are $L_{\alpha}-L_{\beta}$ charge, e.g.
+$Q_{L_{e}-L_{\mu}}=(1,-1,0)$ diagonal matrix.
+The FCNC appears when $V^{\dagger}_{eL(eR)}Q_{L_{\alpha}-L_{\beta}}V_{eL(eR)}$ is not diagonal matrix.
+As an example we consider $M_{e}$ of model (3) in Eq.
+(38) with $(M_{e})_{23}=0$ for simplicity.
+Then, focusing on $e$-$\mu$ sector, the mass matrix is written by
+$M_{e}^{e\mu}=\begin{pmatrix}M_{e}&\delta m\\ 0&M_{\mu}\end{pmatrix}.$ (40)
+Choosing $\delta m\ll M_{\mu}$ this matrix can be approximately diagonalized by
+$V^{\dagger}_{eL}M_{e}^{e\mu}V_{eR}\sim\begin{pmatrix}m_{e}&0\\ 0&m_{\mu}\end{pmatrix},\quad V_{eL}\sim\begin{pmatrix}1&\frac{\delta m}{M_{\mu}}\\ -\frac{\delta m}{M_{\mu}}&1\end{pmatrix},\quad V_{eR}\sim\bf{1},$ (41)
+where $m_{e}\sim M_{e}$ and $m_{\mu}\sim M_{\mu}$. Since model (3) is $U(1)_{L_{e}-L_{\tau}}$ case the LFV interaction is approximately given by
+$g_{X}\frac{\delta m}{m_{\mu}}Z^{\prime}_{\mu}[\bar{\mu}\gamma^{\mu}P_{L}e+\bar{e}\gamma^{\mu}P_{L}\mu].$ (42)
+Then the experimental constraint for $\mu\to ee\bar{e}$ process is most stringent one for heavy $Z^{\prime}$ case  [^3] The case of light $Z^{\prime}$ allowing $\ell_{i}\to\ell_{j}Z^{\prime}$ decay is discussed in ref.
+Asai:2018ocx providing tighter constraints.; that is induced via virtual $Z^{\prime}$ as $\mu\to eZ^{\prime*}\to ee\bar{e}$.
+We can estimate the decay width and branching ratio (BR) as
+$\Gamma_{\mu\to ee\bar{e}}\simeq\frac{g_{X}^{4}}{8m_{Z^{\prime}}^{4}}\left(\frac{\delta m}{m_{\mu}}\right)^{2}\frac{m_{\mu}^{5}}{192\pi^{3}},$ (43)
+$BR(\mu\to ee\bar{e})\simeq\frac{g_{X}^{4}}{g_{2}^{2}G_{F}^{2}m_{Z^{\prime}}^{4}}\left(\frac{\delta m}{m_{\mu}}\right)^{2},$ (44)
+where $G_{F}$ is the Fermi constant. Applying the current limit $BR(\mu\to ee\bar{e})\lesssim 10^{-12}$ SINDRUM:1987nra , we obtain
+$g_{X}^{4}\left(\frac{\delta m}{m_{\mu}}\right)^{2}\left(\frac{\rm TeV}{m_{Z^{\prime}}}\right)^{4}\lesssim 6\times 10^{-11}.$ (45)
+Thus $\delta m/m_{\mu}$ should be very small as $\delta m/m_{\mu}\lesssim\mathcal{O}(10^{-5})$ when $g_{X}=\mathcal{O}(1)$ and $m_{Z^{\prime}}$ is TeV scale.
+Note that we will obtain looser constraint when we consider LFV decay of $\tau$ lepton.
+In the numerical analysis below we assume off-diagonal elements of the matrix is negligibly small and $V_{eL(eR)}\simeq 1$ for simplicity.
 II.4 Neutrino mass
-After spontaneous symmetry breaking we obtain Dirac mass $M_{D}$ between $\nu_{L}$ and $\nu_{R}$ as well as Majorana mass $M_{R}$ of $\nu_{R}$. These mass matrices are written by
+After spontaneous symmetry breaking we obtain Dirac mass $M_{D}$ between $\nu_{L}$ and $\nu_{R}$ ($\overline{\nu_{L}}M_{D}\nu_{R}$) as well as Majorana mass $M_{R}$ of $\nu_{R}$.
+These mass matrices are written by
 $M_{D}$ $=\frac{y_{\nu_{1}}v_{1}}{\sqrt{2}}+\frac{y_{\nu_{2}}v_{2}}{\sqrt{2}},$ (46)
@@
 In Fig.
-1, we show the $\langle m_{ee}\rangle$ and $m_{1}$ (the lightest neutrino mass) for allowed parameter points in each model.
+2, we show the $\langle m_{ee}\rangle$ and $m_{1}$ (the lightest neutrino mass) for allowed parameter points in each model.
 The horizontal lines shows current constraint from KamLand-Zen KamLAND-Zen:2022tow and future prospect in nEXO nEXO:2017nam with energy-density functional (EDF) theory for nuclear matrix element.
@@
 In fig.
-2, we show $\sum m_{\nu}$ and $\delta_{\rm CP}$ for allowed parameter points in each model.
+3, we show $\sum m_{\nu}$ and $\delta_{\rm CP}$ for allowed parameter points in each model.
 The vertical line indicates the upper limit by CMB+DESI data.
@@
 Fig.
-3 and 4 show $\delta_{\rm CP}$-$\alpha_{21}$ and $\delta_{\rm CP}$-$\alpha_{31}$ values for allowed parameter points in each model.
+4 and 5 show $\delta_{\rm CP}$-$\alpha_{21}$ and $\delta_{\rm CP}$-$\alpha_{31}$ values for allowed parameter points in each model.
 We find some correlations among phases where correlations of $\delta_{\rm CP}$-$\alpha_{21}$ in models (1) and (6) [(3) and (5)] provide similar behavior.
@@
 In these plots we show 1(3)$\sigma$ range of $\delta_{\rm CP}$ value by the light-blue(cyan) region as in the Fig.
-2.
+3.
 In addition, we also show correlation between $\alpha_{21}$ and $\alpha_{31}$ in Fig.